OpenStudy (anonymous):

write the rule for each sequence in words. then find the next three terms of 0.7,1.7,2.7,3.7

OpenStudy (anonymous):

add 1.0 0.7+1.0=1.7 1.7+1.0=2.7 2.7+1.0=3.7 3.7+1.0=4.7 and so on ill let you try the final 2 on your own hope this helps

OpenStudy (anonymous):

The rule is "add 1 to any term to get the next term"

OpenStudy (anonymous):

yes